OpenStudy (anonymous):

can anyone help me in this how can i rewrite f(x)=x^2-2x+3 in the form of f(x)=a(x-h)^2+k i need how to do it please include the solution :'( thnx

OpenStudy (amistre64):

complete the square

OpenStudy (amistre64):

we dont 'include the solution' as it doent involve the askers participation.

OpenStudy (anonymous):

i dont know how i will do it -_-"

OpenStudy (amistre64):

completing the square ... im not sure if your lessons have covered it or not, but ultimately that is the process to take

OpenStudy (anonymous):

how i will complete the square in that?

OpenStudy (amistre64):

x^2 - 2x + ____ finding a value that will make this a perfect square another idea, do you know how to find the vertex?

OpenStudy (amistre64):

the point (h,k) defines the vertex and the h value can be determined rather easily

OpenStudy (anonymous):

\[f(x)=x ^{2}-2x+3\] this is the given the problem was i dont know how i can rewrite it

OpenStudy (amistre64):

lets expand and try comparison, since my other ideas dont seem to be ringing any bells ... a(x-h)^2 + k a(x^2 -2hx +h^2) + k ax^2 -2ahx +ah^2 + k given that we need to compare this to: x^2 -2x +3 ax^2 = x^2; what is a? -2ahx = -2x, what is h? ah^2 + k = 3, what is k?